Batch No 465 required $ Kgs of materials in stores which cost $15/Kg, and also a special

component had to be bought at a cost of $1,000. 30 Hours labor were spent on the Batch where the employees were paid at $ 90/Hour. In addition, 3 Hours were spent by supervisor who is paid at $ 150/Hour. Overheads are absorbed at the rate of $ 40/Labor hour. Calculate the total absorption cost of the Batch No 465


1
Expert's answer
2021-11-17T11:10:49-0500

Direct labor costs are $90/hour for 30 hours. Total direct labor costs on the batch are:

90*30 = $2700

Supervisor hours; $150 * 3 = $450

Overheads absorbed at $40/labor hour; thus overhead costs per batch are 40*30 = $1200

Total absorption costs for Batch 465 are:

$15 + $1000 +$2700 + $450 + $1200

= $5365